sir, wat is the difference between argument(z) and amplitude(z) ????please explain..

sir,


wat is the difference between argument(z) and amplitude(z) ????please explain..

Grade:12

1 Answers

Komal
8 years ago
z = a + b i = r /_Ө

r = I z I = [ a² + b² ]

tan Ө = b/a

r is amplitude
Ө is argument